OpenStudy (anonymous):

Find solutions to the equation: xy′+2y=x^3 such that y = 1 when x = 1

OpenStudy (anonymous):

first divide through by x y' + 2 y/x = x^2 you now have the equation in a form where you can apply an integrating factor

OpenStudy (anonymous):

so if I say y = uv and y' = u'v+uv' and put that in I get: u'v+uv' + (2uv)/x = x^2

OpenStudy (anonymous):

Is that what you mean by integrating factor?

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok integrating factor = e ^ INT 2/x dx = e ^ ln x^2 = x^2

OpenStudy (anonymous):

so y IF = INT Q.IF dx where IF = integrating factor and Q = x^2

OpenStudy (anonymous):

x^2 *y = INT x^2 * x^2 dx

OpenStudy (anonymous):

can you carry on from here?

OpenStudy (anonymous):

ok y(x^2) = (x^5)/5 + c

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

now divide by x^2

OpenStudy (anonymous):

y=x^3/5 + c/(x^2)

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

then plug the values in I get 1 = (1^3)/5 + c/(1^2) 1 = 1/5 + c c = 4/5

OpenStudy (anonymous):

y=(x^3)/5+4/(5x^2)

OpenStudy (anonymous):

right the two things to remember with these equations where you cant separate the variables is rearrangge them to the form y' + Py = Q and use y*IF = INT Q*IF dx you can look at it as reverse of Product Rule if u like but for me the easiest way is as I've mentioned

OpenStudy (anonymous):

Thanks a lot for your help

OpenStudy (anonymous):

no probs